OpenStudy (anonymous):

please i need help with algebra. Calculate the slope of the line through your two points and show work. using this equation y-y1=m(x-x1) y is (74,41) and x is (74,42)any help is appreciated im trying to help my cousin to make sure she is doing it right and i have been out of school 20 years now i dont understand her math

OpenStudy (anonymous):

\[y - y _{1} = m(x - x _{1})\] \[m = (y_{2} - y_{1})\div(x_{2} - x_{1})\] \[y_{1} =74, y_{2} = 41\] \[x_{1} = 74, x_{2} = 42\] \[m = (41 - 74) / (42 - 74)\] \[m = 33/32\] The slope is 33/32